I have 3 questions
1.) X to the third + 6xsquared + 10x +0
2.) |5x-4|+2=7
3.)x to the 3/4=-16

Answer by stanbon(75887)

1.) X^3 + 6x^2 + 10x = 0
x(x^2 + 6x + 10) = 0
x = 0 or x^2 + 6x + 10 = 0
x = 0 or x = [-6 +-sqrt(36-4*10)]/2
x = 0 or x = [-6 +- sqrt(-4)]/2
x = 0 or x = -3+i or x = -3-i
---------------------------------------
2.) |5x-4|+2=7
|5x-4| = 5
----
5x -4 = 5 or 5x-4 = -5
5x = 9 or 5x = -1
x = 9/5 or x = -1/5
---------------------------
3.)x to the 3/4=-16
x^(3/4) = -16
---
x = (-16)^(4/3)
----
x = (-8*2)^(4/3
x = (-2*2^(1/3)^4
x = (16*2)2^(1/3)
x = 32*cbrt(2)
-----------------------
Cheers,
Stan H.
-------------